OpenStudy (unique_am_i):

How is this answer correct? 2. A biologist studied the populations of black bear and brown bears over a 10 year period. The biologist modeled the populations, in thousands, with the following polynomials where X is time in years. What polynomial models the total number of brown and black bears? Black bears: 2.3x^2 – 5.6x + 2.3 Brown bears: 2.4x^2 + 7.2x + 0.97 A. [x] 4.7x^2 + 1.6x + 3.27 B. 4.7x^2 – 1.6x – 3.37 C. 4.7x^2 + 1.6x – 3.37 D. 4.7x^2 – 1.6x + 3.37

OpenStudy (baru):

the answer is just the sum of the two polynomials 2.3x^2 +2.4x^2=4.7x^2 -5.6x+7.2x= 1.6x 2.3+0.97=3.27